# Robust Photo-Realistic Hand Gesture Generation: from Single View to Multiple View

### v1

@@ 3.2 Multi-View MANO Mesh Rendering
To maintain proper viewing geometry after the vertex transformation, the camera position must also be adjusted. Specifically, the X-component of the camera translation vector must be mirrored:
$\mathbf{t}_{\text{camera\_rear}}=\begin{pmatrix}-t_{x}\\ t_{y}\\ t_{z}\end{pmatrix},$
where $\mathbf{t}_{\text{camera\_front}}=(t_{x},t_{y},t_{z})^{T}$ is the original camera translation vector in the front view.
@@
where $\mathbf{R}_{x}(\phi)$ is the rotation matrix around the X-axis:
@@
We also adjusted the parameters to make the lighting and colors consistent across the viewpoint meshes. These transformations collectively enable comprehensive multi-view visualization of 3D mesh models while preserving their relative spatial configurations.
3.3 Dual Stream Rendering Encoder for the most Informative Perspective
The information contained in the mesh diagrams of the six viewpoints rendered by the above method is complementary in two ways. Specifically, the front view and the rear view, the left view and the right view, and the top view and the elevation view, respectively, are three sets of complementary viewpoints. Complementary views complement each other’s occluded information. For different gestures, the amount of information contained in these three sets of views is also different.
Based on the characteristics of human gestures, we calculate the projected area of the six views on their respective view planes and select the one with the largest projected area as the most informative view for the gesture. Taking the gesture in Figure 2 as an example, the most informative set of views was calculated to be the front and rear views.
Figure 3: Dual-stream rendering encoder receives a pair of mesh images, extracts the features from them separately, and then fuses the extracted features for output.
@@ 3.3 Dual Stream Rendering Encoder for the most Informative Perspective
Specifically, in each rendering encoder, a pre-trained ResNet50 [6] model, truncated before the final layers, is first used to extract features. Then, a Convolutional Block Attention Module [24] with both channel and spatial attention mechanisms is applied to enhance these features. Finally the feature channel is reduced to 1280 for output by a convolutional layer.
@@
Two rendering encoders are followed by a fusion network, which consists of a series of fully connected layers with ReLU activations [1]. It also employs a residual connection to preserve the original input features while learning new representations.
3.4 Multi-Modal UNet-based Feature Encoder
@@
Figure 4: Training pipeline of Multi-Modal UNet-based Feature Encoder (MUFEN).
@@
Figure 4 demonstrates the training pipeline of Multi-Modal UNet-based Feature Encoder (MUFEN), which consists of two parts: a Multi-Modal Encoder and a Multi-Modal UNet. This architecture is responsible for orchestrating the encoding of different modalities, their fusion, and the subsequent processing through a UNet to produce a final output.
@@
Figure 5: Multi-Modal Encoder integrates features from mesh images, text, depth images, and bounding boxes with a dedicated encoder for each modality.
@@
As shown in Figure 5, the Multi-Modal Encoder encodes and integrates features from different modalities, including mesh images, text, depth images, and bounding boxes. Each modality has a dedicated encoder, and bounding box features are fused into each modality via a tailored fusion mechanism. Specifically, the mesh modality is processed by a Dual Stream Rendering Encoder, which extracts features from two cropped mesh image streams using bounding boxes, as described in the previous section. The depth encoder shares the same structure as a single rendering encoder and extracts both gesture and environmental features from the depth map. Gesture-related text features is extracted from the text label by pre-trained openai-clip-vit-large-patch14. The Bounding Box Encoder built with multi-layer perceptrons (MLP) [9] extracts gesture localization features from the bounding box coordinate data consisting of the 2D coordinates of the top-left and bottom-right corners. These localization features are later used to enhance the spatial localization ability of other modalities through BBox Fusion Modules.
@@
Figure 6: BBox Fusion module adds localization information from the gesture bounding box to the multimodal features extracted by MUFEN through attention and gating mechanisms.
@@
For each modality, there is a dedicated BBox Fusion module, as shown in Figure 6. This module integrates the bounding box features into the modality-specific features using a combination of attention mechanisms and gating strategies. The fusion process involves: Feature Transformation, where the bounding box features are transformed to match the feature dimension of the modality; Attention Fusion, where the bounding box features act as queries, while the modality features serve as keys and values in an attention mechanism, allowing the bounding box features to capture context from the modality features; Learnable Gated Fusion, where a gating mechanism controls the extent to which the bounding box information is integrated into the modality features. By using a learnable gate, the model can adaptively determine the importance of the context features for each token and each sample in the batch, leading to more nuanced and context-aware feature representations; and Residual Connection and Projection, where the final fused features are obtained by adding the original modality features to the gated context features, followed by a linear projection.
@@
The output of the Multi-Modal Encoder is fed into a Multi-Modal UNet. Specifically, modality features fused with bounding box information are first concatenated along the feature dimension and passed through an MLP to obtain a unified fused feature, which then passes through an identity layer, preserving its shape. The spatial dimensions are reduced from 16×16 to 8×8 in the DownBlock, which includes a convolutional layer with self-attention, allowing the model to capture long-range dependencies. At the bottleneck, the features are processed with a self-attention mechanism to further enhance feature interactions without changing the spatial dimensions. During the up-sampling process, the spatial resolution is restored from 8×8 to 16×16 using bilinear interpolation, followed by a convolutional layer. Cross-attention is applied to integrate skip connections from the encoder. Finally, the features are up-scaled to the target resolution of 225×225 and generates the final RGB image output by a convolutional layer.
3.5 Training
@@
Figure 4 demonstrates the training pipeline of MUFEN, which is trained with the following loss function:
$\mathcal{L}=\mathcal{L}_{{\rm{denoise}}}+\lambda\mathcal{L}_{{\rm{reHand}}},$ (4)
@@
where ${{{\mathcal{L}}_{{\rm{reHand}}}}=}{\mathbb{E}}\left|{{\bf I}_{{\rm{reHand}}}-\hat{\bf I}}_{{\rm{reHand}}}\right|$, $\lambda$ is a weight to balance ${{\mathcal{L}}_{\rm{denoise}}}$ and ${{\mathcal{L}}_{\rm{reHand}}}$.
Figure 7: Qualitative Results for different methods. It can be intuitively seen that the hand region generated by the proposed MUFEN is more realistic and more accurately reflect the labels.
4 Experiments
@@
Table 1: Quantitative results. It can be seen that our proposed MUFEN exceeds baselines in all metrics, with substantial improvement in all hand-related metrics. The best quality is highlighted in bold.
@@
SD v1.5 + ControlNet [26] 32.2157 0.0238$\pm$0.0007 31.4976 0.0238$\pm$0.0002
HandRefiner [13] 35.4393 0.0259$\pm$0.0004 35.7291 0.0296$\pm$0.0008
RealisHuman [23] 30.2902 0.0210$\pm$0.0004 31.0369 0.0232$\pm$0.0001
HanDrawer [4] 28.7506 0.0201$\pm$0.0008 26.8279 0.0196$\pm$0.0002
MUFEN (Ours) 26.85 0.0173$\pm$0.0007 26.77 0.0194$\pm$0.0003
4.1 Implementation Details
We used the processed HaGRID v1 dataset provided by  [4], which contained 28,814 training images and 7,623 testing images across 18 gesture categories, with the entire image size is $512\times 512$. Based on the existing multi-modal annotations, we additionally included rear, left, right, top, and bottom views and recorded the corresponding labels and projection areas for each view. For each gesture, we selected the two views with the largest projection areas as its complementary views. During training, we fixed the parameters of the Diffusion v1.5 model and trained the proposed MUFEN together with ControlNet [26] for 90,000 steps using a learning rate of $1{e^{-6}}$ and a batch size of 6. The loss weight $\lambda$ was set to 0.1. For comparison, we used ControlNet [26], HandRefiner [13], RealisHuman [23], and HanDrawer [4] as baselines. All training and inference were conducted on a single NVIDIA Tesla A100 80GB GPU.
4.2 Evaluation Metrics
Frechet Inception Distance (FID) [7] and Kernel Inception Distance (KID) [2] are used for entire image evaluation, and hand region metrics FID-H and KID-H are used to measure the generation quality of hand regions. Hand regions are cropped from an entire image with $299\times 299$ size centred on the hand bounding box, because Inception v3 model for calculating FID requires a $299\times 299$ size image as input. In all cases, lower numerical values represent better performance. FID-H and KID-H are our primary quantitative evaluation metrics as they focus on the hand regions.
4.3 Quantitative Results
@@
The quantitative results comparing several methods are presented in Table 1. Notably, the results indicate that the proposed MUFEN method achieves the best performance across all metrics. For the hand-related metrics, MUFEN achieves an FID-Hand of 26.85 and a KID-Hand of 0.0173±0.0007, showing a substantial improvement compared to the other methods. This comparison indicates that the proposed MUFEN method brings even greater improvements, specifically in the generation of hand regions. This superior performance is primarily attributed to the use of a dual-stream encoder that extracts and fuses complementary information from different viewpoints of the hand. Moreover, the carefully designed multi-modal encoder further enhances the generation capability of the hand modality. These architectural innovations collectively contribute to a more effective representation and synthesis of hand features, as clearly reflected by the improved FID-Hand and KID-Hand metrics.
Figure 8: The performance comparison of different methods on different gestures for entire image. Our proposed MUFEN generally outperforms the baseline methods in terms of FID and KID.
Figure 9: The performance comparison of different methods on different gestures for hand region. Our proposed MUFEN outperforms the baseline methods across most gestures in terms of FID-Hand and KID-Hand.
From Figure 8, we can see two separate comparisons of the five methods—ControlNet, HandRefiner, RealisHuman, HanDrawer, and MUFEN—over 18 hand gesture categories. Regarding the FID plot (left side), MUFEN (in red bars) exhibits the smallest FID in many of the gesture categories, or it is at least very close to the smallest. For example, with the “mute” gesture, MUFEN is noticeably lower than baseline methods. In the middle performance categories like “fist” or “four”, MUFEN again mostly remains below the other methods. At the far right of the FID plot (for example, “rock” gesture), MUFEN still achieves a smaller FID value than other approaches, although all methods have generally higher values there.
Turning to the KID plot (right side in Figure 8), we see MUFEN’s bars remain lower in most categories. For instance, in the “dislike” gesture, MUFEN has the smallest bar, which suggests a significantly better KID compared to the other methods. A similar tendency can be observed in categories such as “one” and “mute,” where MUFEN’s bar is obviously lower than the rest. Although there are some variations among categories—such as “palm” or “fist,” where HanDrawer shows lower KID—the general pattern still indicates MUFEN performs consistently well across the entire set of gestures. We also see that HanDrawer tends to achieve better scores than MUFEN for certain gestures, but it typically does not outperform MUFEN on average.
Figure 9 compares five methods on FID-hand (left chart) and KID-hand (right chart) across eighteen hand gesture categories. Overall, MUFEN achieves consistently low FID-hand and KID-hand values across nearly all hand gestures, indicating superior image fidelity and more accurate hand pose generation. Specifically, in simpler gestures (such as like, dislike, four), MUFEN’s margin of improvement over the next-best method is relatively small because most methods can handle these poses reasonably well. However, as the gestures become more complex – involving multiple specific fingers or unusual orientations (such as inverted peace, mute, OK) – MUFEN’s advantage becomes much more pronounced. In these challenging cases, MUFEN’s errors remain low, while other methods’ errors climb significantly, underscoring MUFEN’s robustness in capturing intricate hand details.
Overall, these bar plots shown in Figure 8 and Figure 9 suggest that MUFEN obtains better performance across most gestures, with especially significant improvements on the most difficult hand poses, indicating that MUFEN’s synthesized hand gestures are more consistent with the real data distribution compared to the other four methods.
4.4 Qualitative Results

### v2

@@ 3.2. Multi-View MANO Mesh Rendering
$\mathbf{t}_{\text{camera\_front}}=(t_{x},t_{y},t_{z})^{T}$ is the original camera translation vector in the front view. To maintain proper viewing geometry after the vertex transformation, the camera position must also be adjusted. Specifically, the X-component of the camera translation vector must be mirrored:
$\mathbf{t}_{\text{camera\_rear}}=\begin{pmatrix}-t_{x}\\ t_{y}\\ t_{z}\end{pmatrix}$.
@@
where $\mathbf{v}_{\text{front}}$ and $\mathbf{v}_{\text{vertical}}$ represent vertex coordinates in the front and vertical viewpoints respectively, and $\mathbf{R}_{x}(\phi)$ is the rotation matrix around the X-axis:
@@
We also adjusted the parameters to make the lighting and colors consistent across the viewpoint meshes. To ensure perceptual uniformity across all viewpoints, the rendering pipeline adopts a unified color parameterization and a standardized multi-light illumination model. The base mesh color is fixed as $\mathbf{C}_{\text{mesh}}=(1.0,1.0,0.9)\in\mathbb{R}^{3}$ with a constant background color $\mathbf{C}_{\text{bg}}=(0,0,0)$, and per-vertex colors are uniformly assigned as $(\mathbf{C}_{\text{mesh}},1.0)$. Left–right hand differentiation is applied via fixed chromatic offsets, ensuring consistent hue mapping across all rendered instances. Material parameters, including metallic factor $m=0.0$ and roughness $r=1.0$, remain fixed to eliminate view-dependent reflectance variations.
Illumination consistency is maintained through a fixed composite lighting model
$\mathbf{L}_{\text{total}}=\mathbf{L}_{\text{ambient}}+\mathbf{L}_{\text{dir}}+\mathbf{L}_{\text{point}}+\mathbf{L}_{\text{raymond}},$
where ambient light $(0.3,0.3,0.3)$ provides a global base tone, and directional, point, and Raymond lights [^1] https://github.com/mmatl/pyrender/blob/master/pyrender/viewer.py are placed at pre-defined positions with constant intensities. View-specific compensation lights are added for oblique or side views while preserving the global intensity balance, ensuring that shading gradients, color perception, and brightness remain consistent across all rendered perspectives.
These transformations collectively enable comprehensive multi-view visualization of 3D mesh models while preserving their relative spatial configurations.
3.3. Dual Stream Rendering Encoder for the most Informative Perspective
The information contained in the six basic view meshes rendered by the above method is complementary and grouped into three complementary pairs: (front, back), (left, right), and (top, bottom). The views in a complementary pair complement each other’s occluded information. For different gestures, the amount of information contained in these three sets of views is also different.
Based on the characteristics of human gestures, we calculate the projected area of the six views on their view planes and select the one with the largest projected area as the most informative view for the gesture. This ensures coverage of critical gesture cues while reducing redundancy. Taking the gesture in Figure 2 as an example, the most informative pair of views was calculated to be the front and rear views.
Figure 3. The dual-stream rendering encoder receives a pair of mesh images, extracts the features from them separately, and then fuses the extracted features for output.
@@ 3.3. Dual Stream Rendering Encoder for the most Informative Perspective
Specifically, in each rendering encoder, a pre-trained ResNet50 (He et al., 2016) model, truncated before the final layers, is first used to extract features. Then, a Convolutional Block Attention Module (Woo et al., 2018) with both channel and spatial attention mechanisms is applied to enhance these features. Finally the feature channel is reduced to 1280 for output by a convolutional layer.
@@
Two rendering encoders are followed by a fusion network, which consists of a series of fully connected layers with ReLU activations (Agarap, 2018). It also employs a residual connection to preserve the original input features while learning new representations.
3.4. Multi-Modal UNet-based Feature Encoder
@@
Figure 4. Training pipeline of Multi-Modal UNet-based Feature Encoder (MUFEN).
@@
Figure 4 demonstrates the training pipeline of Multi-Modal UNet-based Feature Encoder (MUFEN), which consists of two parts: a Multi-Modal Encoder and a Multi-Modal UNet. This architecture is responsible for the encoding of different modalities, their fusion, and the subsequent processing through a UNet to produce a final output.
@@
Figure 5. Our multi-modal encoder integrates features from mesh images, text, depth images, and bounding boxes with a dedicated encoder for each modality.
@@
As shown in Figure 5, the Multi-Modal Encoder encodes and integrates features from different modalities, including mesh images, text, depth images, and bounding boxes. Each modality has a dedicated encoder, and bounding box features are fused into each modality via a tailored fusion mechanism. Specifically, the mesh modality is processed by a Dual Stream Rendering Encoder, which extracts features from two cropped mesh images using bounding boxes, as described in the previous section. The depth encoder shares the same structure as a single rendering encoder and extracts both gesture and environmental features from the depth map. Gesture-related text features are extracted from the text label by pre-trained openai-clip-vit-large-patch14 (Radford et al., 2021). The Bounding Box Encoder built with multi-layer perceptrons (MLP) (Kruse et al., 2022) and extracts gesture localization features from the bounding box coordinate data consisting of the 2D coordinates of the top-left and bottom-right corners. These localization features are later used to enhance the spatial localization ability of other modalities through BBox Fusion modules.
@@
Figure 6. The bounding box (BBox) fusion module adds localization information from the gesture bounding box to the multimodal features extracted by MUFEN through attention and gating mechanisms.
@@
For each modality, there is a dedicated BBox Fusion module, as shown in Figure 6. This module integrates the feature of the bounding box into the feature specific to other modalities using a combination of attention mechanisms and gating strategies. The fusion process involves: Feature Transformation, where the bounding box features are transformed to match the feature dimension of the modality; Attention Fusion, where the bounding box features act as queries, while the modality features serve as keys and values in an attention mechanism, allowing the bounding box features to capture context from the modality features; Learnable Gated Fusion, where a gating mechanism controls the extent to which the bounding box information is integrated into the modality features. By using a learnable gate, the model can adaptively determine the importance of the context features for each token and each sample in the batch, leading to more nuanced and context-aware feature representations; Residual Connection and Projection, where the final fused features are obtained by adding the original modality features to the gated context features, followed by a linear projection.
@@
The output of the Multi-Modal Encoder is fed into a Multi-Modal UNet. Specifically, modality features fused with bounding box information are first concatenated along the feature dimension and passed through an MLP to obtain a unified fused feature, which then passes through an identity layer, preserving its shape. The spatial dimensions are reduced from 16×16 to 8×8 in the DownBlock, which includes a convolutional layer with self-attention, allowing the model to capture long-range dependencies. At the bottleneck, the features are processed with a self-attention mechanism to further enhance feature interactions without changing the spatial dimensions. During the up-sampling process, the spatial resolution is restored from 8×8 to 16×16 using bilinear interpolation, followed by a convolutional layer. Cross-attention is applied to integrate skip connections from the encoder. Finally, the features are up-scaled to the target resolution of 225×225 and generates the final RGB image output using a convolutional layer.
3.5. Training
@@
As shown in Figure 4, MUFEN is trained with the following loss function:
(4) ${\mathcal{L}}=\mathcal{L}_{\mathrm{denoise}}+\lambda{\mathcal{L}}_{{\rm{reHand}}},$
@@
where ${{\mathcal{L}_{\mathrm{reHand}}}=}{\mathbb{E}}\left|{{\bf I}_{{\rm{reHand}}}-\hat{\bf I}}_{{\rm{reHand}}}\right|$, ${\bf I}_{{\rm{reHand}}}$ and ${\hat{\bf I}}_{{\rm{reHand}}}$ are the ground truth gesture region image and the reconstructed gesture region image, respectively, and $\lambda$ is a weight to balance $\mathcal{L}_{\mathrm{denoise}}$} and $\mathcal{L}_{\mathrm{reHand}}$}.
4. Experiments
Figure 7. Qualitative results for different methods.
4.1. Implementation Details
We used the processed HaGRID v1 dataset provided by  (Fu et al., 2025), which contained 28,814 training images and 7,623 testing images across 18 gesture categories, with the entire image size $512\times 512$. Based on the existing multi-modal annotations, we additionally rendered rear, left, right, top, and bottom views mesh using off-the-shelf tool HaMeR (Pavlakos et al., 2024) and recorded the corresponding labels and projection areas for each view. For each gesture, we selected the two views with the largest projection areas as the complementary view pair. During training, we fixed the parameters of the Diffusion v1.5 model and trained the proposed MUFEN together with ControlNet (Zhang et al., 2023) for 90,000 steps using a learning rate of $1{e^{-6}}$ and a batch size of 6. The loss weight $\lambda$ was set to 0.1. For comparison, we used ControlNet (Zhang et al., 2023), HandRefiner (Lu et al., 2024), RealisHuman (Wang et al., 2025), and HanDrawer (Fu et al., 2025) as baselines. All training and inference were conducted on a single NVIDIA Tesla A100 80GB GPU.
4.2. Evaluation Metrics
Frechet Inception Distance (FID) (Heusel et al., 2017) and Kernel Inception Distance (KID) (Bińkowski et al., 2018) are used for entire image evaluation, and hand region metrics FID-Hand and KID-Hand are used to measure the generation quality of hand regions. Hand regions are cropped from an entire image with $299\times 299$ size centred on the hand bounding box, because Inception v3 model for calculating FID requires a $299\times 299$ size image as input. In all cases, lower numerical values represent better performance. FID-H and KID-H are our primary quantitative evaluation metrics as they focus on the hand regions.
@@ 4.2. Evaluation Metrics
Table 1. Quantitative results on HaGRID v1 dataset. MUFEN outperforms existing methods in all metrics, with substantial improvement repetition. The best performance is highlighted in bold.
@@
[…]
Table 2. Paired t-test results of MUFEN compared with other methods over 18 gestures. For each comparison, the $p$-values of both FID-Hand and KID-Hand metrics are reported, along with the number of gestures for which MUFEN achieves better performance.
MUFEN compared With FID-Hand $p$-value Better Count KID-Hand $p$-value Better Count
SD v1.5 + ControlNet (Zhang et al., 2023) $7.6\times 10^{-8}$ 18 $5.6\times 10^{-8}$ 18
HandRefiner (Lu et al., 2024) $3.5\times 10^{-7}$ 18 $3.5\times 10^{-7}$ 18
RealisHuman (Wang et al., 2025) $0.305$ 13 $0.007$ 13
HanDrawer (Fu et al., 2025) $0.059$ 12 $0.018$ 12
Table 3. Ablation results on different modalities and number of views. MUFEN exceeds all ablated variants on hand-region metrics. The best quality is highlighted in bold.
Different settings FID-Hand ↓ KID-Hand ↓ FID ↓ KID ↓
4 views (front, back, left, right) 31.92 0.0222 ± 0.0007 30.52 0.0230 ± 0.0005
4 views (front, back, top, bottom) 32.03 0.0221 ± 0.0007 30.74 0.0235 ± 0.0004
4 views (left, right, top, bottom) 30.78 0.0205 ± 0.0008 28.79 0.0207 ± 0.0004
6 views 39.18 0.0297 ± 0.0009 37.78 0.0308 ± 0.0005
w/o depth map 38.11 0.0298 ± 0.0010 35.37 0.0288 ± 0.0003
w/o mesh 31.37 0.0217 ± 0.0008 31.35 0.0241 ± 0.0003
w/o gesture label 27.88 0.0180 ± 0.0006 27.94 0.0209 ± 0.0001
w/o bbox fusion 27.73 0.0175 ± 0.0007 26.46 0.0186 ± 0.0003
[…]
4.3. Quantitative Results
@@
The quantitative results comparing several methods are presented in Table 1. Notably, the results indicate that the proposed MUFEN method achieves the best performance across all metrics. For the hand-related metrics, MUFEN achieves an FID-Hand of 26.85 and a KID-Hand of 0.0173±0.0007, showing a substantial improvement compared to the other methods. This comparison indicates that the proposed MUFEN method brings great improvements, specifically in the generation of hand regions. This superior performance is primarily attributed to the use of a dual-stream encoder that extracts and fuses complementary information from different viewpoints of the hand. Moreover, the carefully designed multi-modal encoder further enhances the generation capability of the hand modality. These architectural innovations collectively contribute to a more effective representation and synthesis of hand features, as clearly reflected by the improved FID-Hand and KID-Hand metrics.
Table 2 illustrates the paired t-test results of MUFEN compared with other methods over 18 gestures. The paired t-tests across 18 gestures show MUFEN significantly outperforms ControlNet and HandRefiner ($p<10^{-6}$, i.e. better on 18 out of 18 gestures). It also improves over HanDrawer on KID-Hand ($p=0.018$) and is near-significant on FID-Hand ($p=0.059$). Compared to RealisHuman, MUFEN achieves significance on KID-Hand ($p=0.007$), indicating better detail fidelity. These results confirm MUFEN’s advantages over baselines, especially under the KID metric.
To assess computational cost, we compare inference time across single stage inference methods. On a NVIDIA Tesla A100 80GB GPU, MUFEN takes 2.39 seconds to generate an image, slightly faster than HanDrawer (2.45s), marginally slower than MUFEN-single-view (2.12s), and ControlNet is fastest (1.65s). MUFEN offers stronger generation quality with moderate cost. On a NVIDIA GeForce RTX 3080 Laptop GPU, rendering six-view meshes takes 1.45s, compared to 0.48s for front-view-only rendering.
4.4. Qualitative Results